# Trojan Cleansing with Neural Collapse

### v1

@@ 4.1 Neural Collapse Disruption
Table 2: Comparison of cleansing algorithms on trojaned models trained on 10% poisoned data with the BadNet attack. Results on ResNet18 are averaged over 10 random trials with mean±$\pm$±std; result on ViT is a single trial.
@@
Table 3: Comparison of cleansing algorithms on trojaned models trained with 10% poisoned data with the TrojanNN attack. Results on ResNet are averaged over 10 random trials with mean±$\pm$±std; result on ViT is a single trial.
So, while not every metric clearly differentiates trojan and benign models on every dataset, it is the case that benign models always show at least the same degree of convergence as any trojaned model on every collapse metric. And, importantly, on every dataset we observe at least one metric where trojaned models show a significantly weaker degree of collapse than their benign counterpart–suggesting strong tension between Neural Collapse and an effective trojan trigger.
4.2 Trojan Cleansing
We compare our cleansing algorithm, laid out in Section 3.3, with several SOTA cleansing methods on trojaned models. Each of these algorithms uses a small amount of clean data to either analyze or fine-tune the target model. 5% of the training dataset is a common allowance for this setting, but we also run evaluations with 1% of the training dataset as the user in a trojan threat model can be expected to have limited data access. We also use vanilla fine-tuning on the provided dataset as a baseline–these results labeled FT in the following tables.
Our evaluation considers two common trojan attacks: BadNet [10] and TrojanNN [17], and two distinct model architectures: PreAct-ResNet18 [12] and ViT-B/16 [1]. All experiments here were carried out with the BackdoorBench repository [28]. For ResNet, we use the original hyperparameters provided in the configuration. For ViT, we perform moderate hyperparameter tuning for each cleansing method to ensure stable running and reasonable performance. As discussed in Section 3.4, these algorithms may perform better on the transformer models given careful design adaptations, but it is a strength of our algorithm that it readily applies to any classifier architecture out of the box.
@@ 4.2 Trojan Cleansing
BadNet Attack Table 2 presents the cleansing results for the BadNet attack. When the cleansing algorithms are allowed to use 5% of the training data to fine-tune, on the ResNet model, ETF-FT achieves competitive ASR while maintaining high accuracy on clean data, relative to the other methods tested. In more restrictive conditions, where only 1% of the training data is available, ETF-FT and Neural Cleanse maintain strong performance, while other methods show significant degradation.
@@
Vision Transformers Despite their state-of-the-art performance and growing popularity, vision transformers have rarely been examined in the context of backdoor cleansing performance. This is in part due to the fact that their distinct architecture makes it difficult to adapt many common cleansing methods. Table 2 shows cleansing results for the BadNet attack applied to the ViT model, ETF-FT achieves very low ASR in both standard and limited data conditions, surpassing other methods except Neural Cleanse in the standard data case. This does come at the cost of a trade-off with in clean accuracy, which we believe can be mitigated, as discussed in Section 5.1.
@@
TrojanNN Attack Alongside the standard BadNet attack, we consider the more sophisticated TrojanNN attack [17] which identifies specific vulnerable neurons in a pre-trained network and reverse engineers a subtle trigger to target them. As shown in Table 3, ETF-FT is particularly well suited to cleansing this style of trojan. On ResNets, ETF-FT completely eliminates the trigger, achieving an ASR of nearly 0 across 10 trials, while maintaining competitive accuracy with the other cleansing methods. On ViT, many methods, implemented out of the box, fail completely to eliminate the trigger, while ETF-FT achieves an ASR of 10% with only a 1.5 point drop in clean accuracy. The I-BAU algorithm is able to almost entirely eliminate the trigger, but in doing so destroyed the model’s accuracy on clean data.
4.3 Robustness
@@
The above results show that ETF-FT performs well with access to only a small amount of training data. But it is likely that the user of a trojaned model may not have access to any data from the original training distribution at all. We simulated distribution shift in two ways: by imbalancing the class distribution and by applying random erasures within image patches. Results for these datasets, applied to the ResNet BadNet poisoned models, are shown in Tables 5 and 4. In both settings, ETF-FT excels at maintaining the model’s clean accuracy along with competitive ASR relative to other methods. Neural Cleanse, which had generally very strong results in this setting with training data, performs significantly better than all other methods on the imbalanced data, but its clean accuracy suffers on the corrupted data. This is due to the fact that its pruning method relies on identifying important neurons based on their activations on the available data–so it is vulnerable to accidentally pruning neurons associated with valid features if those features aren’t represented in the fine-tuning dataset.
@@
Table 4: Data Corruption: Comparison of cleansing algorithms with 1% corrupted (under random erasure) clean training data on BadNet trojaned ResNet models. Results are averaged over 5 random trials with mean±$\pm$±std.
@@
5 Conclusion
In this paper, we explore how the asymmetry induced by a trojan trigger in a neural network disrupts the phenomenon of Neural Collapse. We provide experimental evidence showing that trojan attacks significantly weaken Neural Collapse based on standard metrics. Leveraging these insights, we propose a novel method for trojan cleansing: by fixing the network’s final layer to a simplex ETF—an ideal structure predicted by Neural Collapse—and fine-tuning on a small amount of clean data. Our results show that this approach effectively removes the trojan trigger while maintaining the network’s accuracy on clean test data without requiring knowledge of the trigger type, the target class, or access to the original training data. Moreover, our method can be applied to any standard architecture with no modification.
5.1 Future Work
This work suggests a variety of interesting future directions. While we’ve shown significant experimental evidence of the connection between trojans and Neural Collapse, it would be valuable to develop a theoretical model of their interaction. Experimentally our results could be extended to different architectures, attack types, and trigger types. In particular, a thorough enough characterization of Neural Collapse in this different settings could allow for the use of our results for detection, not just cleansing, of trojan triggers. And, for the efficacy of our cleansing method, we hope to further explore the choice of the ETF imposed on the network. In this work we used a randomly generated ETF, but, intuitively, an ETF nearer that of the network’s feature means would better maintain clean accuracy, but potentially make it harder to reduce trigger efficacy. We believe careful selection of the ETF there may improve the performance of our cleansing mechanism.

### v2

@@ 4.1 Methodology
Table 2: Comparison of cleansing algorithms on trojaned models trained on 10% poisoned data with the BadNet attack. Results are reported on CIFAR-10. Results on ResNet18 are averaged over 10 random trials with mean±$\pm$±std; result on ViT is a single trial.
Algorithm 1 Trojan Cleansing with Neural Collapse (ETF-FT)
Step 1: Replace weight matrix with $\textbf{W}_{\text{ETF}}$
Construct random ETF $\textbf{W}_{\text{ETF}}$
Replace the final layer weight matrix with $\textbf{W}_{\text{ETF}}$
Step 2: Fine-tune the model
Freeze the final layer weight matrix
Fine-tune on the clean dataset $\mathbf{X}_{\text{clean}}$
Output: Cleansed model $h_{\text{cleansed}}$
ETF-FT is motivated by our results which show that trojans work, in part, by deforming the symmetric structure NC describes in the weights of the final layer. By fixing those weights to be a random, but perfectly symmetric ETF our algorithm accomplishes two things: first, it prevents the trojan deformation of the final layer and second, it resets the weights in a way known to be compatible with good classification. This addresses a challenge present in many trojan cleansing techniques: the more you perturb the poisoned model the less effective the trigger will be, but the harder is it to recover accuracy on clean data.
@@
Table 3: Comparison of cleansing algorithms on trojaned models trained with 10% poisoned data with the TrojanNN attack. Results are reported on CIFAR-10. Results on ResNet are averaged over 10 random trials with mean±$\pm$±std; result on ViT is a single trial.
4.2 Experiment Details
We compare our method with a variety of state-of-the-art cleansing algorithms that have demonstrated strong performance on the BackdoorBench [30] leaderboard. As a baseline, we also run vanilla fine-tuning, with no trojan-aware modifications, on the same subset of clean data–these results are labeled FT in the following tables. Our evaluation considers two types of backdoor attacks: BadNets-A2O [10] and TrojanNN [19], and two model architectures: PreAct-ResNet18 [13] and ViT-B/16 [1].
We focus on two key metrics to evaluate the performance of each cleansing method: accuracy on clean data (ACC) and attack success sate (ASR). Accuracy reflects the model’s performance on clean data after cleansing, ideally showing no reduction compared to the original model. ASR, on the other hand, measures the proportion of poisoned samples classified as the attack target class. An effective cleansing method should reduce ASR to nearly zero.
We also evaluate the cleansing algorithm’s performance on a vision transformer, due to the popularity and efficacy of transformer architectures. However, many cleansing algorithms are not designed with transformers in mind. We adapted algorithms to the new architecture where necessary, as in NPD, which inserts a new layer into the network, we chose to insert a linear layer before the final layer of the transformer encoder. We also did light hyperparameter tuning over batch size, learning rate, scheduler and decay rate for each algorithm including our own. Despite this, many algorithms show poor performance in this setting relative to their results on ResNet. Many of these algorithms may be capable of better performance on transformers given careful adaptation, but in the trojan attack scenario the end user has outsourced network training to an adversary and so is unlikely to be a sophisticated deep learning engineer. As such, it is a significant strength of our algorithm that it can be applied to any classifier architecture without modification.
4.3 Results
@@
BadNet Attack Tab. 2 presents the cleansing results for the BadNet attack. When the cleansing algorithms are allowed to use 5% of the training data to fine-tune, on the ResNet model, ETF-FT achieves competitive ASR while maintaining high accuracy on clean data relative to the other methods tested. In more restrictive conditions, where only 1% of the training data is available, ETF-FT and Neural Cleanse maintain strong performance, while other methods show significant degradation. The numbers in red indicate extremely poor performance, failing to function as an effective cleansing method.
@@
Vision Transformers Despite their state-of-the-art performance and growing popularity, vision transformers have rarely been examined in the context of backdoor cleansing performance. This is in part due to the fact that their distinct architecture makes it difficult to adapt many common cleansing methods. Tab. 2 shows cleansing results for the BadNet attack applied to the ViT model, ETF-FT achieves very low ASR in both standard and limited data conditions, surpassing other methods except Neural Cleanse in the standard data case. This does come at the cost of a trade-off with in clean accuracy, which we believe can be mitigated, as discussed in Sec. 5.1.
@@
TrojanNN Attack Alongside the standard BadNet attack, we consider the more sophisticated TrojanNN attack [19] which identifies specific vulnerable neurons in a pre-trained network and reverse engineers a subtle trigger to target them. As shown in Tab. 3, ETF-FT is particularly well suited to cleansing this style of trojan. On ResNets, ETF-FT completely eliminates the trigger, achieving an ASR of nearly 0 across 10 trials, while maintaining competitive accuracy with the other cleansing methods. On ViT, many methods, implemented out of the box, fail completely to eliminate the trigger, while ETF-FT achieves an ASR of 10% with only a 1.5 point drop in clean accuracy. I-BAU algorithm is able to almost entirely eliminate the trigger, but in doing so destroyed the model’s accuracy on clean data.
4.3.1 Robustness
@@
The above results show that ETF-FT performs well with access to only a small amount of training data. But it is likely that the user of a trojaned model may not have access to any data from the original training distribution at all. We simulated distribution shift in two ways: by imbalancing the class distribution and by applying random erasures within image patches. Results for these datasets, applied to the ResNet BadNet poisoned models, are shown in Tab. 4 and Tab. 8 (in Appendix A). In both settings, ETF-FT excels at maintaining the model’s clean accuracy along with competitive ASR relative to other methods. Neural Cleanse, which had generally very strong results in this setting with training data, performs significantly better than all other methods on the imbalanced data, but its clean accuracy suffers on the corrupted data. This is due to the fact that its pruning method relies on identifying important neurons based on their activations on the available data–so it is vulnerable to accidentally pruning neurons associated with valid features which aren’t represented in the fine-tuning dataset.
@@
Table 4: Data Corruption: Comparison of cleansing algorithms with 1% corrupted (under random erasure) clean training data on BadNet trojaned models. Results are reported on 10% Badnets trojaned CIFAR-10 using PreAct-ResNet18, averaged over 5 random trials with mean±$\pm$±std.
4.3.2 More Sophisticated Trigger
To thoroughly evaluate the performance of ETF-FT against Trojan attacks with more sophisticated triggers beyond the basic yet powerful patch trigger, we conduct experiments using Instagram filter triggers [28]. A demonstration of the filter effects is presented in Fig. 7, where the Gotham filter introduces blurring and a gray tone, while the Lomo filter creates a vignette effect. As shown in Tab. 5, ETF-FT remains effective even against this more subtle type of trigger.
Figure 7: Instagram filter effect: Original (Top), Gotham (Middle), Lomo (Bottom)
Gotham Lomo
Method ACC ASR ACC ASR
Original 93.66 100 93.21 100
FT $89.84\pm 0.35$ $24.5\pm 21.58$ $\textbf{89.92}\pm 0.22$ $3.6\pm 2.11$
NCleanse[29] $84.54\pm 6.05$ ${\color[rgb]{1,0,0}70.49}\pm 28.77$ $83.1\pm 5.18$ $22.68\pm 38.73$
NAD[17] ${\color[rgb]{1,0,0}29.48}\pm 2.61$ $4.14\pm 4.03$ ${\color[rgb]{1,0,0}29.52}\pm 2.48$ $\textbf{0.0}\pm 0$ 0.0
NPD[37] $87.21\pm 2.53$ $\textbf{3.87}\pm 5.29$ $87.68\pm 3.62$ $20.82\pm 18.39$
FT-SAM[36] $85.94\pm 1.42$ $10.59\pm 12.85$ $86.93\pm 1.22$ $1.84\pm 1.03$
I-BAU[33] $\textbf{90.34}\pm 0.86$ $39.12\pm 27.04$ $89.86\pm 1.56$ $8.84\pm 5.58$
ETF-FT(ours) $85.4\pm 0.54$ $14.9\pm 20.7$ $85.33\pm 0.52$ $2.28\pm 1.29$
Table 5: Performance comparison of cleansing algorithms on Trojaned models trained with 10% poisoned data using the BadNets attack, where the trigger is the Gotham and Lomo Instagram filters. Cleansing is performed with 5% clean data. Results are reported on 10% Badnets trojaned CIFAR-10 using PreAct-ResNet18, averaged over 10 random trials (mean $\pm$ std).
4.3.3 Adaptive Attack
Consider a scenario where the adversary is aware that trojan attacks undermine the phenomenon of Neural Collapse. Thus, they train the model with the final layer’s weights fixed as an ETF, effectively replicating the conditions under which Neural Collapse occurs perfectly. ETF-FT demonstrates robust performance even in the face of adaptive attacks. As evidenced in Tab. 6, the ASR is only 90.65, which is lower than the ASR observed in other similar settings. This supports our observation that trojan attacks are incompatible with the Neural Collapse mechanism. Notably, ETF-FT not only removes the trojans but also enhances the model’s Accuracy.
ACC ASR
Original ($W$ fixed as ETF) 87.68 90.65
ETF-FT (5% clean data) $91.48\pm 0.18$ $2.56\pm 0.67$
ETF-FT (1% clean data) $89.91\pm 0.3$ $4.61\pm 4.22$
Table 6: Adaptive Attack: Performance of ETF-FT on trojaned model trained with final layer fixed as ETF. Results are reported on 10% Badnets trojaned CIFAR-10 using PreAct-ResNet18, averaged over 10 random trials (mean $\pm$ std).
4.3.4 Cleansing Without Overtraining
Neural Collapse occurs most strongly in models trained well past zero training error. In line with this, our other experiments used models trained for 350 epochs, despite reaching zero training error at approximately 120 epochs. However, in practical scenarios, models may not always be overtrained. To examine ETF-FT’s reliance on model overtraining we also test it on models trained only until the training error reaches zero. As shown in Tab. 7, ETF-FT maintains strong performance in this case, achieving comparable results to its performance on overtrained models.
ACC ASR
Original 90.66 94.78
ETF-FT (5% clean data) $91.3\pm 0.1$ $0.86\pm 0.23$
ETF-FT (1% clean data) $89.54\pm 0.31$ $0.72\pm 0.23$
Table 7: No Overtraining: Performance of ETF-FT on trojaned model only trained for 120 epochs. Results are reported on 10% Badnets trojaned CIFAR-10 using PreAct-ResNet18, averaged over 10 random trials (mean $\pm$ std).
5 Conclusion
In this paper, we explore how the asymmetry induced by a trojan trigger in a neural network disrupts the phenomenon of Neural Collapse. We provide experimental evidence showing that trojan attacks significantly weaken NC based on standard metrics. Leveraging these insights, we propose a novel method for trojan cleansing: by fixing the network’s final layer to a simplex ETF—an ideal structure predicted by NC—and fine-tuning on a small amount of clean data. Our results show that this approach effectively removes the trojan trigger while maintaining the network’s accuracy on clean test data without requiring knowledge of the trigger type, the target class, or access to the original training data. Moreover, our method can be applied to any standard architecture with no modification.
5.1 Future Work
This work suggests a variety of interesting future directions. Given the significant experimental evidence of the connection between trojans and Neural Collapse, developing a theoretical framework as well as building a trojan detection method ae important areas of future work.
In this work, we used a randomly generated ETF as the starting point in our cleansing strategy; characterizing the impact of the initial ETF selection as well as optimizing this are other interesting areas of investigation.
This paper focuses on image classification, but trojans have also been proven effective in other domains such as language and audio. Extending our analysis and cleansing mechanism to these domains is an interesting direction for future work.
Supplementary Material
Appendix A More Cleansing Results